OpenStudy (anonymous):

The RideEm Bicycles factory can produce 110 bicycles in a day at a total cost of $10,900 and it can produce 130 bicycles in a day at a total cost of $11,600. What are the company's daily fixed costs? What is the marginal cost per bicycle?

OpenStudy (anonymous):

\[y-y_{1}=(\frac{y_{2}-y_{1}}{x_{2}-x_{1}})(x-x_{1})\] \[y-10900=(\frac{11600-10900}{130-110})(x-110)\] \[y-10900=35(x-110)\] \[y-10900=35x-3850\] \[y=35x+7050\]

OpenStudy (anonymous):

$35/bike, fixed cost of 7050

OpenStudy (anonymous):

this is not mathematics this is management. how ever fixed costs mean the Direct labour, direct material and Direct O/H. fixed costs are fixed they dont change per the quantity.

OpenStudy (anonymous):

apply direct variation, then find the constant of proportionality (k)..

OpenStudy (anonymous):

although this is likely wrong... i calculated it using , 110x+C = 10600 130x + C = 11600 20x = 1000 x=50 110(50) + c = 10600 5500+ C = 11600 11600 - 5500 = C